OpenStudy (anonymous):

WILL MEDAL AND FAN!! UST SHOW WORK PLZZ What is the value of 6*7-3^2*9+4^3

OpenStudy (davidusa):

First use PEMDAS. Multiplications first.

OpenStudy (davidusa):

no. exponents.

OpenStudy (davidusa):

What is 3^2 and 4^3? calculate that

OpenStudy (davidusa):

Then multiplication is next.

OpenStudy (anonymous):

3^2= 3*3=9 and 4^3=4*4*4=64 @DavidUsa

OpenStudy (confluxepic):

Parenthesis Exponents Multiplication Division Addition Subtraction

OpenStudy (davidusa):

Yes sir!

OpenStudy (davidusa):

Now do the multiplication.

OpenStudy (davidusa):

here is the new expression 6*7-9*9+64 Do 6*7 and 9*9.

OpenStudy (anonymous):

49-81

OpenStudy (davidusa):

No. 42-81+64

OpenStudy (davidusa):

Now just add that up!

OpenStudy (anonymous):

o lol miss write

OpenStudy (davidusa):

medal and fan? :) do i get one

OpenStudy (anonymous):

yes thx so much

OpenStudy (davidusa):

yay! ur welcome :)